# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ 2.2 Data curation with learning dynamics
We anchor our approach with ideas from learning theory that show the behavior of individual samples during model training (called learning dynamics) contains signal about the nature of the samples themselves (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). Some samples are easily and confidently predicted over different model checkpoints, whereas other samples might be challenging (e.g. due to mislabeling) and hence might be incorrectly predicted for the given label. Consequently, we operationalize learning dynamics as the basis of our curation mechanism. Specifically, we analyze samples in $D_{\mathrm{syn}}$ by studying their learning dynamics computed with a classifier trained on $D_{\mathrm{train}}$. We then categorize and filter samples in $D_{\mathrm{syn}}$, and produce a curated dataset $D_{\mathrm{curated}}\subset D_{\mathrm{syn}}$.
@@
$\bar{P}_{\mathcal{F}}(x,y):=P(\hat{Y}_{\mathcal{F}}(x,y)=1)=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[P(\hat{Y}_{\mathcal{F}}(x,y)=1|H)]=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}$ (1)
@@
$v_{al,\mathcal{F}}(x,y):=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[Var(\hat{Y}_{\mathcal{F}}(x,y)|H)]=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}(1-[f_{e}(x)]_{y})$ (2)
Intuitively, for binary classification ($k=2$), the aleatoric uncertainty for a sample $x$ is maximized when $[f_{e}(x)]_{y}=\frac{1}{2}$ for all checkpoints $f_{e}$, akin to random guessing. Recall aleatoric uncertainty captures the inherent data uncertainty, hence is a principled way to capture issues such as mislabeling. This contrasts epistemic uncertainty, which is model-dependent and can be reduced simply by increasing model parameterization (Hüllermeier and Waegeman, 2021).
Having defined sample-wise confidence and aleatoric uncertainty, we characterize samples in $D_{\mathrm{syn}}$ into two categories, namely Selected and Discarded. Given a sample $(x,y)$, a set of training checkpoints $\mathcal{F}$, and two thresholds $\tau_{\mathrm{conf}}$ and $\tau_{\mathrm{al}}$, we define the category $c(x,y,\mathcal{F})$ as Discarded if $\bar{P}_{\mathcal{F}}(x,y)<\tau_{\mathrm{conf}}$ and $v_{al,\mathcal{F}}(x,y)<\tau_{\mathrm{al}}$, and Selected otherwise.
[…]
To summarize, the objective of the curation step is that training on the curated synthetic data leads to a better classifier $f_{D_{\mathrm{curated}}}$ for the downstream task, compared to training on the uncurated synthetic data, i.e. $M(f_{D_{\mathrm{curated}}})>M(f_{\mathcal{D_{\mathrm{syn}}}})$, where $M$ is a performance measure (for example accuracy). In Sec. 3, we empirically show how performance on this curated dataset is superior both for LLM generated data as well as other classes of generative models.
@@
We highlight two key observations. First, the curated datasets, for all the generative models, exhibit a higher agreement with the proxy $\hat{\eta}$ than the discarded datasets. This aligns with the desideratum of only keeping samples that exhibit the correct feature-label relationships.
@@
This provides a rationale for why curation helps improve discriminative performance, as samples in $D_{\mathrm{curated}}$ are much more likely to have the correct feature-label relationship.
Second, GPT-4 has a higher agreement with $\hat{\eta}$ on $D_{\mathrm{discarded}}$, compared to the other generators. This illustrates that GPT-4’s prior knowledge enables it to better capture the distribution $p(Y|X=x)$. Note that generative baselines (e.g. TVAE) model the joint $p(X,Y)$, without any context of which is the set of features and which is the label. In contrast, we can define in the LLM prompt which column is the target $Y$, allowing the LLM to better capture the feature-label relationships. This complements the findings from Fig. 2, which showed that GPT-4 extrapolates to unseen regions of the feature manifold, captured by the support of $p(X)$.
@@ 3 Curated LLMs for Better Data Augmentation
We now perform an end-to-end quantitative evaluation of CLLM  across multiple real-world datasets, for downstream utility, demonstrating the value of allying the generative capabilities of LLMs with our curation mechanism. Sec. 3.1 compares the performance of GPT-4 and our curation approach with respect to a variety of state-of-the-art tabular augmentation baselines. Having evaluated CLLM  on a range of datasets, we also demonstrate how we can leverage information extracted during curation to characterize datasets via a hardness proxy. Sec. 3.2 illustrates how our characterization of samples during the curation step can help to flag synthesized datasets (e.g via the LLM) which, if used for training, will result in poor downstream performance.
Experimental setup. We compare CLLM  (with GPT-4 (OpenAI, 2023) and GPT-3.5 (Brown et al., 2020)) against a variety of baselines for tabular data generation and augmentation: CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), TabDDPM (Kotelnikov et al., 2022), SMOTE (Chawla et al., 2002) and GReaT (Borisov et al., 2023), which fine-tunes an LLM. We evaluate performance on 7 real-world datasets with different feature counts and vary the number of samples available in $D_{\mathrm{train}}$, repeating each experiment across $10$ seeds.
While we do not know the exact makeup of the pretraining data of LLMs like GPT-4, there is the possibility that open-source data might be included. This poses a risk of memorization as the primary source of performance gain. To disentangle the role of memorization, we select 4 real-world medical datasets (Maggic (Pocock et al., 2013), Covid (Baqui et al., 2020), SEER (Duggan et al., 2016), CUTRACT (PCUK, 2019)) that require an authorization process to access, hence are unlikely to form part of the LLMs training corpus. We use common open-source datasets (Adult and Drug from the UCI repository (Asuncion and Newman, 2007) and Compas (Angwin et al., 2016)) that are highly reflective of data scarce domains. Further experimental details can be found in Appendix B.
@@
covid (n=20) 74.41 68.50 73.78 73.87 normal-$\uparrow$ 69.85 71.41 $\uparrow$ 59.00 63.67 $\uparrow$ 66.84 66.85 $\uparrow$ 57.38 66.46 $\uparrow$ 62.87 68.56 $\uparrow$ 66.95 66.82 61.69 66.11 $\uparrow$
cutract (n=20) 72.23 70.12 71.15 72.50 normal-$\uparrow$ 69.97 71.54 $\uparrow$ 64.01 67.98 $\uparrow$ 66.05 66.59 $\uparrow$ 52.38 67.02 $\uparrow$ 64.44 70.42 $\uparrow$ 68.41 69.24 $\uparrow$ 68.94 70.22 $\uparrow$
maggic (n=20) 67.41 57.13 60.70 61.48 normal-$\uparrow$ 57.54 58.69 $\uparrow$ 52.75 54.51 $\uparrow$ 54.59 55.39 $\uparrow$ 50.29 55.64 $\uparrow$ 54.72 57.38 $\uparrow$ 55.84 56.15 $\uparrow$ 54.08 56.19 $\uparrow$
seer (n=20) 87.92 80.67 84.53 84.82 normal-$\uparrow$ 83.34 83.71 $\uparrow$ 74.34 78.73 $\uparrow$ 80.59 80.60 $\uparrow$ 47.57 74.43 $\uparrow$ 76.06 79.98 $\uparrow$ 79.23 80.02 $\uparrow$ 74.53 78.73 $\uparrow$
@@
drug (n=20) 77.81 70.84 75.08 75.29 normal-$\uparrow$ 71.68 72.14 $\uparrow$ 68.31 72.65 $\uparrow$ 68.12 69.68 $\uparrow$ 58.78 68.89 $\uparrow$ 62.13 67.75 $\uparrow$ 70.16 70.16 66.60 69.18 $\uparrow$
covid (n=40) 75.02 70.77 73.40 73.95 normal-$\uparrow$ 70.42 71.93 $\uparrow$ 63.63 68.46 $\uparrow$ 70.50 70.44 56.50 68.68 $\uparrow$ 66.41 70.48 $\uparrow$ 68.66 68.44 61.03 67.35 $\uparrow$
cutract (n=40) 72.57 69.18 69.87 71.72 normal-$\uparrow$ 68.47 69.56 $\uparrow$ 63.01 67.87 $\uparrow$ 65.63 67.27 $\uparrow$ 54.39 68.44 $\uparrow$ 61.40 67.98 $\uparrow$ 67.86 67.95 $\uparrow$ 59.79 66.62 $\uparrow$
maggic (n=40) 67.50 58.26 59.29 60.77 normal-$\uparrow$ 57.50 59.15 $\uparrow$ 55.00 56.78 $\uparrow$ 55.24 56.94 $\uparrow$ 48.81 56.64 $\uparrow$ 54.68 58.58 $\uparrow$ 57.40 57.44 $\uparrow$ 55.04 57.33 $\uparrow$
seer (n=40) 87.90 82.93 84.29 84.93 normal-$\uparrow$ 83.46 84.44 $\uparrow$ 80.05 83.67 $\uparrow$ 82.59 81.37 54.93 81.11 $\uparrow$ 79.88 84.36 $\uparrow$ 80.79 82.21 $\uparrow$ 78.69 83.62 $\uparrow$
compas (n=40) 67.35 62.34 67.57 67.85 normal-$\uparrow$ 61.34 62.84 $\uparrow$ 56.29 61.02 $\uparrow$ 58.85 60.11 $\uparrow$ 58.88 64.37 $\uparrow$ 58.61 63.54 $\uparrow$ 60.83 60.95 $\uparrow$ 55.94 61.04 $\uparrow$
adult (n=40) 84.43 79.44 48.31 73.82 $\uparrow$ 49.21 74.27 $\uparrow$ 71.82 79.11 $\uparrow$ 71.51 77.99 $\uparrow$ 66.77 78.81 $\uparrow$ 71.13 79.71 $\uparrow$ 77.90 78.84 $\uparrow$ 72.58 80.02 normal-$\uparrow$
drug (n=40) 77.71 71.86 74.30 75.79 normal-$\uparrow$ 71.33 72.76 $\uparrow$ 69.46 72.74 $\uparrow$ 71.08 73.07 $\uparrow$ 64.89 73.64 $\uparrow$ 62.51 70.97 $\uparrow$ 69.23 69.78 $\uparrow$ 65.22 70.30 $\uparrow$
covid (n=100) 74.52 71.57 73.77 74.71 normal-$\uparrow$ 70.71 72.76 $\uparrow$ 69.05 72.13 $\uparrow$ 71.60 73.22 $\uparrow$ 63.52 72.04 $\uparrow$ 64.25 72.64 $\uparrow$ 70.08 70.78 $\uparrow$ 69.05 71.96 $\uparrow$
cutract (n=100) 72.36 70.96 70.20 72.51 normal-$\uparrow$ 69.97 71.94 $\uparrow$ 67.94 72.42 $\uparrow$ 70.53 71.98 $\uparrow$ 55.72 69.14 $\uparrow$ 67.59 72.42 $\uparrow$ 68.79 69.68 $\uparrow$ 66.89 71.52 $\uparrow$
maggic (n=100) 67.46 59.65 58.98 61.32 normal-$\uparrow$ 55.71 58.90 $\uparrow$ 57.20 59.34 $\uparrow$ 57.26 58.28 $\uparrow$ 49.54 57.91 $\uparrow$ 56.36 60.11 $\uparrow$ 58.89 58.99 $\uparrow$ 56.17 58.86 $\uparrow$
seer (n=100) 87.79 83.95 84.45 85.37 normal-$\uparrow$ 83.92 85.08 $\uparrow$ 81.60 85.14 $\uparrow$ 83.04 84.83 $\uparrow$ 70.32 83.83 $\uparrow$ 81.16 85.03 $\uparrow$ 81.82 82.49 $\uparrow$ 78.88 84.50 $\uparrow$
compas (n=100) 67.18 62.56 68.02 68.19 normal-$\uparrow$ 60.10 62.47 $\uparrow$ 60.01 63.73 $\uparrow$ 58.32 61.34 $\uparrow$ 59.97 64.19 $\uparrow$ 60.02 64.04 $\uparrow$ 61.44 61.73 $\uparrow$ 59.97 62.82 $\uparrow$
adult (n=100) 84.34 81.24 46.09 74.57 $\uparrow$ 47.56 73.97 $\uparrow$ 74.29 80.45 $\uparrow$ 75.93 78.22 $\uparrow$ 77.09 81.66 normal-$\uparrow$ 70.70 81.04 $\uparrow$ 80.56 81.10 $\uparrow$ 74.04 80.23 $\uparrow$
drug (n=100) 78.00 73.58 76.24 76.74 normal-$\uparrow$ 69.46 71.05 $\uparrow$ 68.19 73.28 $\uparrow$ 72.43 73.79 $\uparrow$ 67.26 75.28 $\uparrow$ 62.67 73.12 $\uparrow$ 70.90 71.53 $\uparrow$ 68.22 73.59 $\uparrow$
covid (n=200) 74.69 72.33 73.40 74.62 normal-$\uparrow$ 70.70 73.12 $\uparrow$ 71.07 73.89 $\uparrow$ 72.47 74.44 $\uparrow$ 65.55 73.07 $\uparrow$ 65.04 72.90 $\uparrow$ 71.68 71.87 $\uparrow$ 67.89 72.38 $\uparrow$
cutract (n=200) 72.52 71.75 71.39 73.01 $\uparrow$ 70.28 72.39 $\uparrow$ 69.28 72.41 $\uparrow$ 71.83 74.03 normal-$\uparrow$ 66.66 72.49 $\uparrow$ 68.77 73.16 $\uparrow$ 70.23 70.80 $\uparrow$ 66.61 71.87 $\uparrow$
maggic (n=200) 67.37 61.39 58.92 61.41 normal-$\uparrow$ 57.33 60.16 $\uparrow$ 58.48 61.33 $\uparrow$ 56.26 57.20 $\uparrow$ 50.74 59.60 $\uparrow$ 55.95 60.75 $\uparrow$ 60.73 60.78 $\uparrow$ 57.18 60.23 $\uparrow$
seer (n=200) 87.84 84.63 84.39 85.56 $\uparrow$ 83.48 84.80 $\uparrow$ 82.04 85.34 $\uparrow$ 84.39 86.57 normal-$\uparrow$ 82.15 86.03 $\uparrow$ 77.73 85.19 $\uparrow$ 83.38 84.15 $\uparrow$ 79.71 85.26 $\uparrow$
compas (n=200) 67.14 63.27 67.02 68.15 normal-$\uparrow$ 60.48 63.39 $\uparrow$ 60.58 64.32 $\uparrow$ 60.60 63.52 $\uparrow$ 61.11 65.08 $\uparrow$ 56.58 63.60 $\uparrow$ 61.99 62.80 $\uparrow$ 60.15 63.99 $\uparrow$
adult (n=200) 84.25 82.12 40.96 75.84 $\uparrow$ 49.89 76.11 $\uparrow$ 78.18 82.32 $\uparrow$ 81.66 83.17 $\uparrow$ 80.06 83.32 normal-$\uparrow$ 74.31 82.64 $\uparrow$ 82.26 82.39 $\uparrow$ 75.21 82.02 $\uparrow$
drug (n=200) 77.36 76.10 75.58 76.06 $\uparrow$ 70.66 72.81 $\uparrow$ 71.31 75.98 $\uparrow$ 69.61 71.79 $\uparrow$ 72.35 77.41 normal-$\uparrow$ 65.25 75.26 $\uparrow$ 74.38 74.78 $\uparrow$ 68.39 74.33 $\uparrow$

### v3

@@ 2.2 Data curation with learning dynamics
We anchor our approach with ideas from learning theory that show that the behavior of individual samples during model training (called learning dynamics) contains signal about the nature of the samples themselves (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). Some samples are easily and confidently predicted over different model checkpoints, whereas other samples might be challenging (e.g. due to mislabeling) and hence might be incorrectly predicted for the given label. Consequently, we operationalize learning dynamics as the basis of our proposed curation mechanism. Specifically, we analyze samples in $D_{\mathrm{syn}}$ by studying their learning dynamics computed with a classifier trained on $D_{\mathrm{train}}$. We then categorize and filter samples in $D_{\mathrm{syn}}$, and produce a curated dataset $D_{\mathrm{curated}}\subset D_{\mathrm{syn}}$.
@@
$\bar{P}_{\mathcal{F}}(x,y)$ $:=P(\hat{Y}_{\mathcal{F}}(x,y)=1)$
$=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[P(\hat{Y}_{\mathcal{F}}(x,y)=1|H)]$
$=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}$
@@
$v_{al,\mathcal{F}}(x,y)$ $:=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[Var(\hat{Y}_{\mathcal{F}}(x,y)|H)]$
$=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}(1-[f_{e}(x)]_{y})$
Intuitively, for binary classification ($k=2$), the aleatoric uncertainty for a sample $x$ is maximized when $[f_{e}(x)]_{y}=\frac{1}{2}$ for all checkpoints $f_{e}$, akin to random guessing. Recall aleatoric uncertainty captures the inherent data uncertainty, hence is a principled way to capture issues such as mislabeling. This contrasts epistemic uncertainty, which is model-dependent and can be reduced simply by increasing model parameterization (Hüllermeier & Waegeman, 2021).
Having defined sample-wise confidence and aleatoric uncertainty, we categorize samples in $D_{\mathrm{syn}}$ as Selected or Discarded: for a sample $(x,y)$, a set of training checkpoints $\mathcal{F}$, and two thresholds $\tau_{\mathrm{conf}}$ and $\tau_{\mathrm{al}}$, we define the category $c(x,y,\mathcal{F})$ as Discarded if $\bar{P}_{\mathcal{F}}(x,y)<\tau_{\mathrm{conf}}$ and $v_{al,\mathcal{F}}(x,y)<\tau_{\mathrm{al}}$, and Selected otherwise.
Hence, a Discarded sample is one for which we have a very low confidence in predicting its associated label whereas we also have low inherent data uncertainty. Finally, given a function $f$ associated with the set of checkpoints $\mathcal{F}$, we define the curated set $D_{\mathrm{curated}}=\{(x,y)|(x,y)\in D_{\mathrm{syn}},c(x,y,\mathcal{F})=\textit{Selected}\}$. We also define $D_{\mathrm{discarded}}=D_{\mathrm{syn}}\setminus D_{\mathrm{curated}}$.
To summarize, the objective of the curation step is that training on the curated synthetic data leads to a better classifier $f_{D_{\mathrm{curated}}}$ for the downstream task, compared to training on the uncurated synthetic data, i.e. $M(f_{D_{\mathrm{curated}}})>M(f_{\mathcal{D_{\mathrm{syn}}}})$, where $M$ is a performance measure (for example accuracy). In Sec. 3, we empirically show how performance on this curated dataset is superior both for LLM generated data, as well as other classes of generative models.
@@
We highlight two key observations. First, the curated datasets, for all the generative models, exhibit a higher agreement with the proxy $\hat{\eta}$ than the discarded datasets. This aligns with the desideratum of only keeping samples that exhibit the correct feature-label relationships. This provides a rationale for why curation helps improve discriminative performance, as samples in $D_{\mathrm{curated}}$ are much more likely to have the correct feature-label relationship.
Second, GPT-4 has a higher agreement with $\hat{\eta}$ on $D_{\mathrm{discarded}}$, compared to other generators. This illustrates that GPT-4’s prior knowledge enables it to better capture the distribution $p(Y|X=x)$. Note that generative baselines (e.g. TVAE) model the joint $p(X,Y)$, without any context of which is the set of features and which is the label. In contrast, we can define in the LLM prompt which column is the target $Y$, allowing the LLM to better capture the feature-label relationships. This complements the findings from Fig. 2, which showed that GPT-4 extrapolates to unseen regions of the feature manifold, captured by the support of $p(X)$.
@@ 3 Curated LLMs for Better Data Augmentation
We now perform an end-to-end quantitative evaluation of CLLM   [^2] Code: https://github.com/seedatnabeel/CLLM or https://github.com/vanderschaarlab/CLLM across multiple real-world datasets, for downstream utility, demonstrating the value of allying the generative capabilities of LLMs with our curation mechanism.
Sec. 3.1 compares the downstream performance of models when trained on uncurated vs curated data for a variety of augmentation approaches. Having evaluated CLLM  on a range of datasets, we also demonstrate how we can leverage information extracted during curation to characterize datasets via a hardness proxy. Sec. 3.2 illustrates how our characterization of samples during the curation step can help to flag synthesized datasets (e.g via the LLM) which, if used for training, will result in poor downstream performance.
Table 3: AUC averaged over 4 downstream models on $D_{\mathrm{test}}$. Curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$. CLLM  w/ GPT-4 (Cur.) provides the strongest performance for both private/proprietary datasets and public datasets
@@
covid (n=20) 74.41 68.50 73.78 73.87 $\uparrow$ 69.85 71.41 $\uparrow$ 59.00 63.67 $\uparrow$ 66.84 66.85 $\uparrow$ 57.38 66.46 $\uparrow$ 62.87 68.56 $\uparrow$ 66.95 66.82 61.69 66.11 $\uparrow$
cutract (n=20) 72.23 70.12 71.15 72.50 $\uparrow$ 69.97 71.54 $\uparrow$ 64.01 67.98 $\uparrow$ 66.05 66.59 $\uparrow$ 52.38 67.02 $\uparrow$ 64.44 70.42 $\uparrow$ 68.41 69.24 $\uparrow$ 68.94 70.22 $\uparrow$
maggic (n=20) 67.41 57.13 60.70 61.48 $\uparrow$ 57.54 58.69 $\uparrow$ 52.75 54.51 $\uparrow$ 54.59 55.39 $\uparrow$ 50.29 55.64 $\uparrow$ 54.72 57.38 $\uparrow$ 55.84 56.15 $\uparrow$ 54.08 56.19 $\uparrow$
seer (n=20) 87.92 80.67 84.53 84.82 $\uparrow$ 83.34 83.71 $\uparrow$ 74.34 78.73 $\uparrow$ 80.59 80.60 $\uparrow$ 47.57 74.43 $\uparrow$ 76.06 79.98 $\uparrow$ 79.23 80.02 $\uparrow$ 74.53 78.73 $\uparrow$
@@
drug (n=20) 77.81 70.84 75.08 75.29 $\uparrow$ 71.68 72.14 $\uparrow$ 68.31 72.65 $\uparrow$ 68.12 69.68 $\uparrow$ 58.78 68.89 $\uparrow$ 62.13 67.75 $\uparrow$ 70.16 70.16 66.60 69.18 $\uparrow$
covid (n=40) 74.41 70.77 73.40 73.95 $\uparrow$ 70.42 71.93 $\uparrow$ 63.63 68.46 $\uparrow$ 70.50 70.44 56.50 68.68 $\uparrow$ 66.41 70.48 $\uparrow$ 68.66 68.44 61.03 67.35 $\uparrow$
cutract (n=40) 72.23 69.18 69.87 71.72 $\uparrow$ 68.47 69.56 $\uparrow$ 63.01 67.87 $\uparrow$ 65.63 67.27 $\uparrow$ 54.39 68.44 $\uparrow$ 61.40 67.98 $\uparrow$ 67.86 67.95 $\uparrow$ 59.79 66.62 $\uparrow$
maggic (n=40) 67.41 58.26 59.29 60.77 $\uparrow$ 57.50 59.15 $\uparrow$ 55.00 56.78 $\uparrow$ 55.24 56.94 $\uparrow$ 48.81 56.64 $\uparrow$ 54.68 58.58 $\uparrow$ 57.40 57.44 $\uparrow$ 55.04 57.33 $\uparrow$
seer (n=40) 87.92 82.93 84.29 84.93 $\uparrow$ 83.46 84.44 $\uparrow$ 80.05 83.67 $\uparrow$ 82.59 81.37 54.93 81.11 $\uparrow$ 79.88 84.36 $\uparrow$ 80.79 82.21 $\uparrow$ 78.69 83.62 $\uparrow$
compas (n=40) 67.51 62.34 67.57 67.85 $\uparrow$ 61.34 62.84 $\uparrow$ 56.29 61.02 $\uparrow$ 58.85 60.11 $\uparrow$ 58.88 64.37 $\uparrow$ 58.61 63.54 $\uparrow$ 60.83 60.95 $\uparrow$ 55.94 61.04 $\uparrow$
adult (n=40) 84.17 79.44 48.31 73.82 $\uparrow$ 49.21 74.27 $\uparrow$ 71.82 79.11 $\uparrow$ 71.51 77.99 $\uparrow$ 66.77 78.81 $\uparrow$ 71.13 79.71 $\uparrow$ 77.90 78.84 $\uparrow$ 72.58 80.02 $\uparrow$
drug (n=40) 77.81 71.86 74.30 75.79 $\uparrow$ 71.33 72.76 $\uparrow$ 69.46 72.74 $\uparrow$ 71.08 73.07 $\uparrow$ 64.89 73.64 $\uparrow$ 62.51 70.97 $\uparrow$ 69.23 69.78 $\uparrow$ 65.22 70.30 $\uparrow$
covid (n=100) 74.41 71.57 73.77 74.71 $\uparrow$ 70.71 72.76 $\uparrow$ 69.05 72.13 $\uparrow$ 71.60 73.22 $\uparrow$ 63.52 72.04 $\uparrow$ 64.25 72.64 $\uparrow$ 70.08 70.78 $\uparrow$ 69.05 71.96 $\uparrow$
cutract (n=100) 72.23 70.96 70.20 72.51 $\uparrow$ 69.97 71.94 $\uparrow$ 67.94 72.42 $\uparrow$ 70.53 71.98 $\uparrow$ 55.72 69.14 $\uparrow$ 67.59 72.42 $\uparrow$ 68.79 69.68 $\uparrow$ 66.89 71.52 $\uparrow$
maggic (n=100) 67.41 59.65 58.98 61.32 $\uparrow$ 55.71 58.90 $\uparrow$ 57.20 59.34 $\uparrow$ 57.26 58.28 $\uparrow$ 49.54 57.91 $\uparrow$ 56.36 60.11 $\uparrow$ 58.89 58.99 $\uparrow$ 56.17 58.86 $\uparrow$
seer (n=100) 87.92 83.95 84.45 85.37 $\uparrow$ 83.92 85.08 $\uparrow$ 81.60 85.14 $\uparrow$ 83.04 84.83 $\uparrow$ 70.32 83.83 $\uparrow$ 81.16 85.03 $\uparrow$ 81.82 82.49 $\uparrow$ 78.88 84.50 $\uparrow$
compas (n=100) 67.51 62.56 68.02 68.19 $\uparrow$ 60.10 62.47 $\uparrow$ 60.01 63.73 $\uparrow$ 58.32 61.34 $\uparrow$ 59.97 64.19 $\uparrow$ 60.02 64.04 $\uparrow$ 61.44 61.73 $\uparrow$ 59.97 62.82 $\uparrow$
adult (n=100) 84.17 81.24 46.09 74.57 $\uparrow$ 47.56 73.97 $\uparrow$ 74.29 80.45 $\uparrow$ 75.93 78.22 $\uparrow$ 77.09 81.66 $\uparrow$ 70.70 81.04 $\uparrow$ 80.56 81.10 $\uparrow$ 74.04 80.23 $\uparrow$
drug (n=100) 77.81 73.58 76.24 76.74 $\uparrow$ 69.46 71.05 $\uparrow$ 68.19 73.28 $\uparrow$ 72.43 73.79 $\uparrow$ 67.26 75.28 $\uparrow$ 62.67 73.12 $\uparrow$ 70.90 71.53 $\uparrow$ 68.22 73.59 $\uparrow$
covid (n=200) 74.41 72.33 73.40 74.62 $\uparrow$ 70.70 73.12 $\uparrow$ 71.07 73.89 $\uparrow$ 72.47 74.44 $\uparrow$ 65.55 73.07 $\uparrow$ 65.04 72.90 $\uparrow$ 71.68 71.87 $\uparrow$ 67.89 72.38 $\uparrow$
cutract (n=200) 72.23 71.75 71.39 73.01 $\uparrow$ 70.28 72.39 $\uparrow$ 69.28 72.41 $\uparrow$ 71.83 74.03 $\uparrow$ 66.66 72.49 $\uparrow$ 68.77 73.16 $\uparrow$ 70.23 70.80 $\uparrow$ 66.61 71.87 $\uparrow$
maggic (n=200) 67.41 61.39 58.92 61.41 $\uparrow$ 57.33 60.16 $\uparrow$ 58.48 61.33 $\uparrow$ 56.26 57.20 $\uparrow$ 50.74 59.60 $\uparrow$ 55.95 60.75 $\uparrow$ 60.73 60.78 $\uparrow$ 57.18 60.23 $\uparrow$
seer (n=200) 87.92 84.63 84.39 85.56 $\uparrow$ 83.48 84.80 $\uparrow$ 82.04 85.34 $\uparrow$ 84.39 86.57 $\uparrow$ 82.15 86.03 $\uparrow$ 77.73 85.19 $\uparrow$ 83.38 84.15 $\uparrow$ 79.71 85.26 $\uparrow$
compas (n=200) 67.51 63.27 67.02 68.15 $\uparrow$ 60.48 63.39 $\uparrow$ 60.58 64.32 $\uparrow$ 60.60 63.52 $\uparrow$ 61.11 65.08 $\uparrow$ 56.58 63.60 $\uparrow$ 61.99 62.80 $\uparrow$ 60.15 63.99 $\uparrow$
adult (n=200) 84.17 82.12 40.96 75.84 $\uparrow$ 49.89 76.11 $\uparrow$ 78.18 82.32 $\uparrow$ 81.66 83.17 $\uparrow$ 80.06 83.32 $\uparrow$ 74.31 82.64 $\uparrow$ 82.26 82.39 $\uparrow$ 75.21 82.02 $\uparrow$
drug (n=200) 77.81 76.10 75.58 76.06 $\uparrow$ 70.66 72.81 $\uparrow$ 71.31 75.98 $\uparrow$ 69.61 71.79 $\uparrow$ 72.35 77.41 $\uparrow$ 65.25 75.26 $\uparrow$ 74.38 74.78 $\uparrow$ 68.39 74.33 $\uparrow$
Experimental setup. We compare CLLM  (with GPT-4 (OpenAI, 2023) and GPT-3.5 (Brown et al., 2020)) against a variety of baselines for tabular data generation and augmentation: CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), TabDDPM (Kotelnikov et al., 2022), SMOTE (Chawla et al., 2002) and GReaT (Borisov et al., 2023), which fine-tunes an LLM. We evaluate performance on 7 real-world datasets with different feature counts and representative of the diverse domains where CLLM  can have impact. For each dataset, we vary the number of samples available in $D_{\mathrm{train}}$, repeating each experiment for $10$ seeds.
While we do not know the exact makeup of the pretraining data for LLMs like GPT-4, there is the possibility that open-source data might be included. This poses the risk of memorization as the primary source of performance gain. To disentangle the role of memorization, we select 4 real-world medical datasets (Maggic (Pocock et al., 2013), Covid (Baqui et al., 2020), SEER (Duggan et al., 2016), CUTRACT (PCUK, 2019)) that require an authorization process to access, hence are unlikely to form part of the LLMs training corpus. We use common open-source datasets (Adult and Drug from the UCI repository (Asuncion & Newman, 2007) and Compas (Angwin et al., 2016)) that are highly reflective of data scarce domains. Further experimental details can be found in Appendix B.